### 26 Mar 2026 v1

# Highly Efficient Rank-Adaptive Sweep-based SI-DSA for the Radiative Transfer Equation via Mild Space Augmentation

Abstract
Low-rank methods have emerged as a promising strategy for reducing the memory footprint and computational cost of discrete-ordinates discretizations of the radiative transfer equation (RTE). However, most existing rank-adaptive approaches rely on rank-proportional space augmentation, which can negate efficiency gains when the effective solution rank becomes moderately large. To overcome this limitation, we develop a rank-adaptive sweep-based source iteration with diffusion synthetic acceleration (SI–DSA) for the first-order steady-state RTE. The core of our method is a sweep-based inner-loop iterative low-rank solver that performs efficient rank adaptation via mild space augmentation. In each inner iteration, the spatial basis is augmented with a small, rank-independent number of basis vectors without truncation, while a single truncation is performed only after the inner loop converges. Efficient rank adaptation is achieved through residual-based greedy angular subsampling strategy together with incremental updates of projection operators, enabling non-intrusive reuse of existing transport-sweep implementations. In the outer iteration, a DSA preconditioner is applied to accelerate convergence. Numerical experiments show that the proposed solver achieves accuracy and iteration counts comparable to those of full-rank SI–DSA while substantially reducing memory usage and runtime, even for challenging multiscale problems in which the effective rank reaches $30\text{–}45\%$ of the full rank.
1 Introduction
The radiative transfer equation (RTE) is a fundamental kinetic equation describing the transport and interaction of particles, such as photons and neutrons, as they propagate through and interact with a background medium. It has a wide range of applications in medical imaging, astrophysics and nuclear engineering. The unknown of RTE, commonly referred to as the angular flux or particle distribution depending on the underlying applications, is defined over a high-dimensional phase space comprising the spatial locations and angular directions of particle transport. This high dimensionality renders deterministic discretizations computationally demanding in both memory and runtime.
To mitigate the computational challenges, low-rank methods have been developed to approximate solutions of both steady-state and time-dependent RTE via matrix or tensor factorizations. Existing approaches for time-dependent RTE mainly follow two paradigms: dynamical low-rank (DLR) methods [koch2007dynamical], which evolve the solution directly on a low-rank manifold by projecting the underlying dynamical system onto the tangent space; and step-and-truncation (SAT) strategies [dektor2021rank, guo2022low], which first evolve the solution in an enlarged space and then project it back onto a low-rank manifold via truncation or rounding. Instances of low-rank method for time dependent RTE include DLR methods for the first- or second-order formulations of RTE [peng2020low, peng2023sweep, haut2026efficient], DLR and SAT method based on micro-macro decompositions [einkemmer2021asymptotic, einkemmer2024asymptotic, frank2025asymptotic, sands2024high], high-order low-order (HOLO) DLR method [peng2021high], and rank-adaptive predictor-corrector approaches [hauck2023predictor]; these have been further analyzed in [ding2021dynamical, yin2025towards]. For steady state RTE, preconditioned rank-adaptive iterative solvers have been developed, including soft-thresholding based low-rank Richardson iteration with an exponential sum preconditioner [bachmayr2024low] and a low-rank source iteration with diffusion synthetic acceleration (SI-DSA) based on the second order reformulation of RTE [guo2025inexact]. For a systematic review of low-rank methods solving kinetic equations, we refer to [einkemmer2025review]. For a broader overview on low-rank methods for general PDEs, see [bachmayr2023low]. Despite these advances in low-rank methodologies, achieving substantial computational gains is often neither straightforward nor guaranteed, particularly for challenging practical problems whose solutions do not exhibit extremely low-rank structure. The main reason is that some intermediate steps can be computationally expensive (e.g., scaling superlinearly in the solution rank) and may offset the efficiency gains of the low-rank approach. Consequently, the design of low-rank schemes that deliver genuine speedups beyond idealized test cases remains challenging.
In this work, we focus on the steady-state RTE, for which preconditioned iterative methods, such as source iteration (SI) and Krylov methods with diffusion synthetic acceleration (DSA) preconditioners, are well established [Adams2002FastIM]. Although low-rank iterative solvers for the RTE remain comparatively underdeveloped, several pioneering efforts have emerged in recent years. Existing low-rank iterative methods can be broadly classified into two main categories. The first is based on a second-order reformulation of the RTE, within which low-rank methods have been developed under both rank-adaptive and DLR frameworks. Representative examples include the rank-adaptive preconditioned low-rank Richardson iteration in [bachmayr2017iterative], the rank-adaptive SI-DSA in [guo2025inexact], and the fixed-rank DLR approach in [haut2026efficient]. The second category directly targets the first-order formulation of the RTE and exploits transport sweeps, a fast matrix-free algorithm to solve the discrete streaming system. Existing methods along this direction include the projection-based fixed-rank DLR method in [peng2023sweep] and the collocation-based fixed-rank $S_{N}$-like DLR method in [haut2026efficient]. Despite these encouraging developments, several limitations remain:
  1. 1.
    Ill-posedness of the second-order formulation: When the total cross section is small or vanishes in a sub-region of the computational domain, the linear system resulting from the second-order reformulation of RTE becomes ill-posed or even singular, posing computational challenges to methods built upon it [bachmayr2024low, guo2025inexact, haut2026efficient].
  2. 2.
    Fixed-rank constraints in sweep-based solvers: Current sweep-based methods for the first-order formulation in [peng2023sweep, haut2026efficient] are restricted to fixed-rank approximations. However, the optimal rank required to balance accuracy and efficiency is typically unknown a priori. To the best of our knowledge, rank adaptivity has not yet been integrated into sweep-based low-rank iterative solvers for the first-order RTE.
  3. 3.
    Overhead of rank-adaptive space augmentation: Standard rank-adaptation strategies often rely on $r$-dependent space augmentation, where $r$ denotes the solution rank. Typically, these methods enrich the mode space by increasing its dimension to a multiple of $r$ before truncation (e.g., $2r$ in augmented basis update Galerkin (BUG) [ceruti2022rank] or $4r$ for the low-rank SI-DSA in [guo2025inexact] when solving 2D2V RTE). However, such augmentation can be overly aggressive for challenging problems that require a relatively high solution rank to achieve a prescribed accuracy or satisfy the convergence criterion of the iterative solver. For example, the effective rank may reach approximately $40\%$ of the full rank [guo2025inexact] in order to reduce the difference between the scalar fluxes of successive iterations below $10^{-5}$. In these regimes, the computational overhead of $r$-dependent augmentation can offset performance gains, preventing genuine computational savings compared to efficient full-rank solvers.
To address these challenges, we develop an efficient rank-adaptive low-rank SI-DSA based on a mild augmentation procedure. Specifically, we design a novel rank-adaptation strategy that avoids rank-dependent, overly aggressive basis enrichment. In our method, the enrichment is performed progressively, and the dimension of the resulting enriched mode space is only slightly larger than $r$ prior to truncation, rather than a multiple thereof. The key components of our approach are as follows.
  1. 1.
    A novel iterative low-rank solver for the SI step is designed to achieve efficient rank-adaptation in an incremental manner. During each inner-loop iteration, a small, constant number of modes is added without immediate truncation; a single truncation is then performed upon convergence of the inner loop. This is realized through a novel residual-based greedy random subsampling of angular directions.
  2. 2.
    To further enhance efficiency of the low-rank SI method, the spatial low-rank basis and the projected operators are updated incrementally, thereby avoiding unnecessary global orthogonalizations and projections.
  3. 3.
    Within the SI-DSA framework, our low-rank SI serves as a seamless alternative to the full-rank SI, enabling the direct application of DSA and transport sweeps.
The primary contribution of this paper is the development of, to the best of our knowledge, the first rank-adaptive, sweep-based iterative solver for the first order formulation of the RTE. Central to our approach is a novel, efficient rank-adaptive technique that requires only mild space augmentation, ensuring computational speedups even for challenging problems with moderate-to-high solution ranks, e.g., $40\%$ of the full rank.
We note that our method shares conceptual similarities with the fixed-rank, collocation-based DLR method recently proposed in the concurrent work [haut2026efficient]. Both frameworks construct low-rank approximations by combining angular space sampling with projections in the physical space. However, beyond the capability for rank adaptivity, our approach differs fundamentally in the algorithmic integration of the angular sampling. The method in [haut2026efficient] is designed for implicit time-marching: at each time step, a fixed number of angles is sampled based on the solution from the previous step, and a modified discrete ordinate system determined by these angles is subsequently solved via SI-DSA. Consequently, their sampling occurs externally to the iterative solver. In contrast, our method adaptively samples angles to update the solution rank internally within the inner-loop iterations of the low-rank SI. This integrated design enables our rank adaptation through mild, rather than aggressive, space augmentation.
Our paper is organized as follows. In Sec. 2, we introduce the model problem and motivate the development of our method. In Sec. 3, we develop our low-rank SI-DSA. In Sec. 4, we numerically demonstrate the performance of our method through a series of benchmark tests. At last, we draw our conclusions in Sec. 5.
2 Background
In this paper, we consider steady-state, single group, linear RTE with isotropic scattering and isotropic inflow boundary conditions on the computational domain $\Gamma_{\mathbf{x}}$:
(1a) $\bm{\Omega}\cdot\nabla\psi(\mathbf{x},\bm{\Omega})+\sigma_{t}(\mathbf{x})\psi(\mathbf{x},\bm{\Omega})=\sigma_{s}(\mathbf{x})\phi(\mathbf{x})+G(\mathbf{x}),\quad\sigma_{t}(\mathbf{x})=\sigma_{s}(\mathbf{x})+\sigma_{a}(\mathbf{x}),$
(1b) $\phi(\mathbf{x})=\frac{1}{4\pi}\int_{\mathbb{S}^{2}}\psi(\mathbf{x},\bm{\Omega})d\bm{\Omega},\quad\mathbf{x}\in\Gamma_{\mathbf{x}},\;\bm{\Omega}\in\mathbb{S}^{2},$
(1c) $\psi(\mathbf{x},\bm{\Omega})=0,\quad\mathbf{x}\in\partial\Gamma_{\mathbf{x}},\;\bm{\Omega}\cdot\bm{n}(\mathbf{x})<0.$
The RTE is posed in the $(\bm{\Omega},\mathbf{x})$-phase space, where $\bm{\Omega}=(\bm{\Omega}_{x},\bm{\Omega}_{y},\bm{\Omega}_{z})\in\mathbb{S}^{2}$ and $\mathbf{x}=(x,y,z)\in\Gamma_{\mathbf{x}}$ denote the angular direction and spatial location, respectively. The angular flux at location $\mathbf{x}$ in direction $\bm{\Omega}$ is denoted by $\psi(\mathbf{x},\bm{\Omega})$, while the scalar flux is defined as $\phi(\mathbf{x})=\frac{1}{4\pi}\int_{\mathbb{S}^{2}}\psi(\mathbf{x},\bm{\Omega})d\bm{\Omega}$. $G(\mathbf{x})$ represents an isotropic source. The scattering, absorption and total cross sections of the background medium are denoted by $\sigma_{s}(\mathbf{x})\geq 0$, $\sigma_{a}(\mathbf{x})\geq 0$ and $\sigma_{t}(\mathbf{x})=\sigma_{s}(\mathbf{x})+\sigma_{a}(\mathbf{x})$, respectively. The inflow boundary condition is given in (1c), where $\bm{n}(\mathbf{x})$ denotes the outward normal direction of the computational domain at location $\mathbf{x}\in\partial\Gamma_{\mathbf{x}}$.
As the scattering cross section $\sigma_{s}\rightarrow\infty$, the solution of the RTE (1) formally converges to its diffusion limit: $\psi(\mathbf{x},\bm{\Omega})\rightarrow\phi(\mathbf{x})$, where $\phi(\mathbf{x})$ solves the diffusion equation
(2) $\nabla\cdot\left(\frac{1}{\sigma_{s}}\bm{\mathcal{D}}\nabla\phi\right)=-\sigma_{a}\phi+G\;\text{with}\;\bm{\mathcal{D}}=\frac{1}{4\pi}\textrm{diag}\left(\int_{\mathbb{S}^{2}}\bm{\Omega}_{x}^{2}d\bm{\Omega},\int_{\mathbb{S}^{2}}\bm{\Omega}_{y}^{2}d\bm{\Omega},\int_{\mathbb{S}^{2}}\bm{\Omega}_{z}^{2}d\bm{\Omega}\right).$
2.1 $S_{N}$ upwind discontinuous Galerkin discretization
Following [Adams2002FastIM, larsen2009advances], we employ a discrete ordinates $(S_{N})$ angular discretization [pomraning2005equations] coupled with an upwind discontinuous Galerkin (DG) spatial discretization. As shown in [adams2001discontinuous, guermond2010asymptotic], this specific discretization is asymptotic preserving when utilizing piecewise linear or higher order polynomials. Consequently, the diffusion limit can be captured without requiring a refined mesh to resolve the particle mean free path. The details of our numerical scheme are presented below.
$S_{N}$ angular discretization: We solve RTE (1) at a set of quadrature points $\{\bm{\Omega}_{j}\}_{j=1}^{N_{\bm{\Omega}}}$ in the angular space and approximate the scalar flux through numerical quadrature. In particular, we seek $\psi(\mathbf{x},\bm{\Omega}_{j})\approx\psi_{j}(\mathbf{x})$ for $j=1,\dots,N_{\bm{\Omega}}$,
(3a) $\bm{\Omega}_{j}\cdot\nabla\psi_{j}(\mathbf{x})+\sigma_{t}\psi(\mathbf{x})=\sigma_{s}\phi(\mathbf{x})+G(\mathbf{x}),\quad\phi(\mathbf{x})=\sum_{j=1}^{N_{\bm{\Omega}}}\omega_{j}\psi_{j}(\mathbf{x}),\;\mathbf{x}\in\Gamma_{\mathbf{x}}$
(3b) $\psi_{j}(\mathbf{x})=g(\mathbf{x}),\quad\mathbf{x}\in\partial\Gamma_{\mathbf{x}},\bm{\Omega}_{j}\cdot\bm{n}(x)<0,$
where $\{\omega_{j}\}_{j=1}^{N_{\bm{\Omega}}}$ denote the normalized quadrature weights associated with $\{\bm{\Omega}_{j}\}_{j=1}^{N_{\bm{\Omega}}}$ satisfying $\sum_{j=1}^{N_{\bm{\Omega}}}\omega_{j}=1$. In this paper, we employ the Chebyshev–Legendre quadrature; see Appx. A for details.
Upwind DG method to solve the $S_{N}$ system (3): For simplicity, we focus on a 2D rectangular spatial domain $\Gamma_{x}=[x_{L},x_{R}]\times[y_{B},y_{T}]$ partitioned by a Cartesian mesh $\mathcal{T}_{h}=\{\mathcal{T}_{a,b}=[x_{a-\frac{1}{2}},x_{a+\frac{1}{2}}]\times[y_{b-\frac{1}{2}},y_{b+\frac{1}{2}}],\;1\leq a\leq N_{x},\;1\leq b\leq N_{y}\}$ with $x_{a-\frac{1}{2}}=x_{L}+(a-\frac{1}{2})\frac{x_{R}-x_{L}}{N_{x}}$ and $y_{b-\frac{1}{2}}=y_{B}+(b-\frac{1}{2})\frac{y_{T}-y_{B}}{N_{y}}$. Specifically, $x_{\frac{1}{2}}=x_{L}$, $x_{N_{x}+\frac{1}{2}}=x_{R}$, $y_{\frac{1}{2}}=y_{B}$ and $y_{N_{y}+\frac{1}{2}}=y_{T}$. For $j=1,2,\dots,N_{\bm{\Omega}}$, we seek the numerical solution $\psi(\mathbf{x},\bm{\Omega}_{j})\approx\psi_{h,j}\in U_{h}^{K}(\mathcal{T}_{h})$ in the discrete space
(4a) $U_{h}^{K}(\mathcal{T}_{h}):=\{u(\mathbf{x}):u(\mathbf{x})|_{\mathcal{T}_{i}}\in Q^{K}(\mathcal{T}_{a,b}),1\leq a\leq N_{x},1\leq b\leq N_{y}\},$
(4b) $Q^{K}(\mathcal{T}_{a,b})=\{x^{p}y^{q},0\leq p,q\leq K,(x,y)\in\mathcal{T}_{a,b}\}.$
Here, $Q^{K}(\mathcal{T}_{a,b})$ denotes the space of tensor-product polynomials of degree at most $K$ in each coordinate on the cell $\mathcal{T}_{a,b}$. More specifically, the functions $\psi_{h,j}$ satisfy $\forall\eta_{h}(\mathbf{x})\in U_{h}^{K}(\mathcal{T}_{h}),$
$\sum_{a=1}^{N_{x}}\sum_{b=1}^{N_{y}}\Bigg(-\int_{\mathcal{T}_{a,b}}\Big(\bm{\Omega}_{j}\cdot\nabla\eta_{h}\Big)\psi_{h,j}d\mathbf{x}+\int_{\mathcal{T}_{a,b}}\sigma_{t}\psi_{h,j}\eta_{h}d\mathbf{x}$
$+\int_{y_{b-\frac{1}{2}}}^{y_{b+\frac{1}{2}}}\bm{\Omega}_{j,x}\widehat{\psi}_{h,j}(x_{a+\frac{1}{2}},y)\eta_{h}(x_{a+\frac{1}{2}}^{-},y)dy-\int_{y_{b-\frac{1}{2}}}^{y_{b+\frac{1}{2}}}\bm{\Omega}_{j,x}\widehat{\psi}_{h,j}(x_{a+\frac{1}{2}},y)\eta_{h}(x_{a+\frac{1}{2}}^{+},y)dy$
$+\int_{x_{a-\frac{1}{2}}}^{y_{a+\frac{1}{2}}}\bm{\Omega}_{j,y}\widehat{\psi}_{h,j}(x,y_{b+\frac{1}{2}})\eta_{h}(x,y_{b+\frac{1}{2}}^{-})dx-\int_{x_{a-\frac{1}{2}}}^{x_{a+\frac{1}{2}}}\bm{\Omega}_{j,y}\widehat{\psi}_{h,j}(x,y_{b-\frac{1}{2}})\eta_{h}(x,y_{b-\frac{1}{2}}^{+})dx\Bigg)$
$=\sum_{a=1}^{N_{x}}\sum_{b=1}^{N_{y}}\left(\int_{\mathcal{T}_{a,b}}\sigma_{s}(\mathbf{x})\phi_{h}(\mathbf{x})\eta_{h}(\mathbf{x})d\mathbf{x}+\int_{\mathcal{T}_{a,b}}G(\mathbf{x})\eta_{h}(\mathbf{x})d\mathbf{x}\right),$
where $\phi_{h}=\sum_{j=1}^{N_{\bm{\Omega}}}\omega_{j}\psi_{h,j}$, $h(\alpha^{\pm},y)=\lim_{x\rightarrow\alpha^{\pm}}h(\alpha,y)$, $h(x,\beta^{\pm})=\lim_{y\rightarrow\beta^{\pm}}h(x,y)$ and the upwind flux $\widehat{\psi}_{h,j}$ is defined as
(5a) $\bm{\Omega}_{j,x}\widehat{\psi}_{h,j}(x_{a-\frac{1}{2}},y)=\bm{\Omega}_{j,x}\psi_{h,j}(x_{a-\frac{1}{2}}^{\mp},y),\quad\text{if }\pm\bm{\Omega}_{j,x}\geq 0,\;1\leq a\leq N_{x}+1,$
(5b) $\bm{\Omega}_{j,y}\widehat{\psi}_{h,j}(x,y_{b-\frac{1}{2}})=\bm{\Omega}_{j,y}\psi_{h,j}(x,y_{b-\frac{1}{2}}^{\mp}),\quad\text{if }\pm\bm{\Omega}_{j,y}\geq 0,\;1\leq b\leq N_{y}+1.$
In addition, the boundary fluxes are determined by $\psi_{h,j}(x_{\frac{1}{2}}^{-},y)=g(x_{L},y)$ for $\bm{\Omega}_{j,x}\geq 0$, $\psi_{h,j}(x_{N_{x}+\frac{1}{2}}^{+},y)=g(x_{R},y)$ for $\bm{\Omega}_{j,x}\leq 0$, $\psi_{h,j}(x,y_{\frac{1}{2}}^{-})=g(x,y_{B})$ for $\bm{\Omega}_{j,y}\geq 0$ and $\psi_{h,j}(x,y_{N_{y}+\frac{1}{2}}^{+})=g(x,y_{T})$ for $\bm{\Omega}_{j,y}\leq 0$.
Matrix-vector formulation of the fully discrete system. Let $\{\eta_{k}\}_{i=1}^{N_{\mathbf{x}}}$ with $N_{\mathbf{x}}=N_{x}N_{y}(K+1)^{2}$ be an orthonormal basis for $U_{h}^{K}$, and assume that $\eta_{i}$ is locally supported in the element $\mathcal{T}_{a_{i},b_{i}}$, i.e., $\eta_{i}(x,y)=0$ if $(x,y)\not\in\mathcal{T}_{a_{i},b_{i}}$. The numerical solution $\psi_{h,j}$ can be expanded as $\psi_{h,j}=\sum_{i=1}^{N_{\mathbf{x}}}\psi_{ij}$. Let $\bm{\psi}_{j}=(\psi_{1j},\dots,\psi_{N_{\mathbf{x}}j})^{T}\in\mathbb{R}^{N_{\mathbf{x}}}$, which satisfies the following linear system:
(6a) $({\bf{D}}_{j}+{\mathbf{\Sigma}}_{t})\bm{\psi}_{j}={\mathbf{\Sigma}}_{s}\bm{\phi}+\bm{G}+\bm{g}_{j}^{\textrm{bc}}={\mathbf{\Sigma}}_{s}\bm{\phi}+\widetilde{\bm{G}}_{j},\quad\bm{\phi}=\sum_{j=1}^{N_{\bm{\Omega}}}\omega_{j}\bm{\psi}_{j},$
(6b) where ${\bf{D}}_{j}=\begin{cases}\bm{\Omega}_{j,x}{\bf{D}}_{x}^{-}+\bm{\Omega}_{j,y}{\bf{D}}_{y}^{-},\;\text{if }\bm{\Omega}_{j,x}\geq 0,\bm{\Omega}_{j,y}\geq 0,\\ \bm{\Omega}_{j,x}{\bf{D}}_{x}^{-}+\bm{\Omega}_{j,y}{\bf{D}}_{y}^{+},\;\text{if }\bm{\Omega}_{j,x}\geq 0,\bm{\Omega}_{j,y}<0,\\ \bm{\Omega}_{j,x}{\bf{D}}_{x}^{+}+\bm{\Omega}_{j,y}{\bf{D}}_{y}^{-},\;\text{if }\bm{\Omega}_{j,x}<0,\bm{\Omega}_{j,y}\geq 0,\\ \bm{\Omega}_{j,x}{\bf{D}}_{x}^{+}+\bm{\Omega}_{j,y}{\bf{D}}_{y}^{+},\;\text{otherwise}.\end{cases}$
Here, ${\bf{D}}_{j}\in\mathbb{R}^{N_{\mathbf{x}}\times N_{\mathbf{x}}}$ denotes the discrete transport operator associated with $\bm{\Omega}_{j}\cdot\nabla$, ${\mathbf{\Sigma}}_{s},{\mathbf{\Sigma}}_{t}\in\mathbb{R}^{N_{\mathbf{x}}\times N_{\mathbf{x}}}$ denote the discrete scattering and total cross-section operators, and $\bm{G},\,\bm{g}_{j}^{\textrm{bc}}\in\mathbb{R}^{N_{\mathbf{x}}}$ correspond to the discretizations of the source term and the boundary conditions, respectively. For brevity, the precise definitions of ${\bf{D}}_{x}^{\pm},{\bf{D}}_{y}^{\pm},{\mathbf{\Sigma}}_{t},{\mathbf{\Sigma}}_{s},\bm{G},$ and $\bm{g}_{j}^{\mathrm{bc}}$ are provided in Appx. B.
2.2 Transport sweep and source iteration with diffusion synthetic acceleration (SI-DSA)
SI-DSA [Adams2002FastIM, larsen2009advances] is a popular preconditioned iterative method to solve (6). A crucial component of the SI framework is the transport sweep [Adams2002FastIM, larsen2009advances], a highly efficient, matrix-free algorithm designed to invert the discrete streaming operator ${\bf{D}}_{j}+{\mathbf{\Sigma}}_{t}$. Below, we briefly review the fundamentals of SI-DSA and the associated matrix-free transport sweeps.
SI-DSA. Given an initial guess for the scalar flux $\bm{\phi}^{(0)}$, the $n$-th iteration of SI-DSA follows a two-step procedure.
  1. 1.
    The SI step updating the angular flux: We update the angular flux by solving the transport equations for each discrete ordinate independently, using the scalar flux from the previous iteration:
    (7) $({\bf{D}}_{j}+{\mathbf{\Sigma}}_{t})\bm{\psi}_{j}^{(n)}={\mathbf{\Sigma}}_{s}\bm{\phi}^{(n-1)}+\widetilde{\bm{G}}_{j},$
    and then update the scalar flux as $\phi^{(n,*)}=\sum_{j=1}^{N_{\bm{\Omega}}}\omega_{j}\bm{\psi}_{j}^{(n)}$.
  2. 2.
    The DSA step accelerating convergence. DSA introduces a correction to the scalar flux by solving:
    (8) ${\bf{D}}_{\textrm{diff.}}\delta\bm{\phi}^{(n)}=-{\mathbf{\Sigma}}_{a}\delta\bm{\phi}^{(n)}+{\mathbf{\Sigma}}_{s}(\bm{\phi}^{(n,*)}-\bm{\phi}^{(n)}),$
    and then corrects scalar flux as $\bm{\phi}^{(n)}=\bm{\phi}^{(n,*)}+\delta\bm{\phi}^{(n)}$. Equation (8) represents the discretization of the diffusion limit for the error equation
    (9) $\frac{1}{3}\nabla\cdot\left(\frac{1}{\sigma_{s}}\nabla\delta\phi^{(n)}\right)=-\sigma_{a}\delta\bm{\phi}^{(n)}+\sigma_{s}(\phi^{(n,*)}-\phi^{(n)}).$
    We emphasize that, as discussed in [Adams2002FastIM, larsen2009advances, palii2020convergent], a consistent discretization of (9) is required to ensure effective acceleration. Specifically, we employ a fully consistent discretization following [adams2001discontinuous], with detailed definitions provided in Appx. B of [peng2024reduced]. We also note that, as shown in [larsen2009advances], SI may converge arbitrarily slowly without DSA.
The convergence criteria of SI-DSA is typically chosen as $\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{\infty}\leq\epsilon_{\textrm{SI-SA}}$, where $\epsilon_{\textrm{SI-SA}}$ is a prescribed threshold.
Transport sweeps. Transport sweeps leverage the following property to solve (7) very efficiently: with upwind discretization, the discrete streaming operator, ${\bf{D}}_{j}+{\mathbf{\Sigma}}_{t}$, becomes block lower triangular when elements are ordered along the upwind direction corresponding to $\bm{\Omega}_{j}$. Hence, the SI equation (7) can be solved very efficiently in a matrix-free manner through a single block Gauss–Seidel iteration sweeping elements along the upwind direction.
To better illustrate the transport sweep mechanism, we describe the block structure and sweeping procedure for the RTE in 1D slab geometry:
(10) $v\partial_{x}\psi+\sigma_{t}\psi=\sigma_{s}\phi+G,\;x\in[x_{L},x_{R}],v\in[-1,1]$
with inflow boundary conditions. We partition the computational domain as
$x_{L},x_{R}]=\cup_{i=1}^{N_{x}}I_{a},\;I_{a}=[x_{L}+(a-1)\Delta x,x_{R}+a\Delta x],\;\Delta x=(x_{R}-x_{L})/N_{x}$
and apply $S_{N}$ upwind DG discretization using piecewise polynomials of degree at most $K$. The resulting linear system retains the form (6a). Define $\bm{\psi}^{(n)}_{j,a}\in\mathbb{R}^{K+1}$ as the vector of DOFs of $\bm{\psi}_{j}^{(n)}$ associated with local basis functions in $I_{a}$. By ordering the elements along the upwind direction corresponding to $v_{j}$, namely from $1$ to $N_{x}$ for $v_{j}>0$ and from $N_{x}$ to $1$ for $v_{j}<0$, the SI equation takes a block lower triangular form:
(11a) $\left(\begin{matrix}\mathbf{A}^{[j]}_{1,1}&\\ \mathbf{A}^{[j]}_{2,1}&\mathbf{A}^{(v_{j})}_{2,2}&\\ &\ddots&\ddots&\\ &&\mathbf{A}^{[j]}_{N_{x},N_{x}-1}&\mathbf{A}^{[j]}_{N_{x},N_{x}}\end{matrix}\right)\left(\begin{matrix}\bm{\psi}^{(n)}_{j,1}\\ \bm{\psi}^{(n)}_{j,2}\\ \vdots\\ \bm{\psi}^{(n)}_{j,N_{x}}\end{matrix}\right)=\left(\begin{matrix}\mathbf{b}^{(n)}_{j,1}\\ \mathbf{b}^{(n)}_{j,2}\\ \vdots\\ \mathbf{b}^{(n)}_{j,N_{x}}\end{matrix}\right),\;\text{with }v_{j}\geq 0,$
(11b) $\left(\begin{matrix}\mathbf{A}^{[j]}_{N_{x},N_{x}}&\\ \mathbf{A}^{[j]}_{N_{x}-1,N_{x}}&\mathbf{A}^{[j]}_{N_{x}-1,N_{x}-1}&\\ &\ddots&\ddots&\\ &&\mathbf{A}^{[j]}_{1,2}&\mathbf{A}^{[j]}_{1,1}\end{matrix}\right)\left(\begin{matrix}\bm{\psi}^{(n)}_{j,N_{x}}\\ \bm{\psi}^{(n)}_{j,N_{x}-1}\\ \vdots\\ \bm{\psi}^{(n)}_{j,1}\end{matrix}\right)=\left(\begin{matrix}\mathbf{b}^{(n)}_{j,N_{x}}\\ \mathbf{b}^{(n)}_{j,N_{x}-1}\\ \vdots\\ \mathbf{b}^{(n)}_{j,1}\end{matrix}\right),\;\text{with }v_{j}<0.$
As a result, $\bm{\psi}_{j}^{(n)}$ can be computed through a single transport sweep, which is equivalent to a single block Gauss-Seidel iteration:
  • For $v_{j}\geq 0$, sweep from $a=1$ to $a=N_{x}$ and solve $\mathbf{A}_{a,a}^{[j]}\bm{\psi}_{j,a}^{(n)}=\mathbf{b}^{(n)}_{j,a}-\mathbf{A}_{a,a-1}^{[j]}\bm{\psi}_{j,a-1}$ with $\bm{\psi}_{j,0}^{(n)}=\mathbf{0}$.
  • For $v_{j}<0$, sweep from $a=N_{x}$ to $a=1$ and solve $\mathbf{A}_{a,a}^{[j]}\bm{\psi}_{j,a}^{(n)}=\mathbf{b}^{(n)}_{j,a}-\mathbf{A}_{a,a+1}^{[j]}\bm{\psi}_{j,a+1}$ with $\bm{\psi}_{j,N_{x}+1}^{(n)}=\mathbf{0}$.
This sweeping procedure can be straightforwardly extended to a 2D rectangular computational domain partitioned by a Cartesian mesh.
2.3 Low-rank solvers for steady-state RTE and challenges
To mitigate the substantial memory and computational costs arising from the high dimensionality of the phase space, low-rank methods approximate the discrete solution using low-rank matrix or tensor decompositions. For example,
(12) $\begin{bmatrix}\bm{\psi}(\cdot,\bm{\Omega}_{1})&\bm{\psi}(\cdot,\bm{\Omega}_{2})&\dots&\bm{\psi}(\cdot,\bm{\Omega}_{N_{\bm{\Omega}}})\end{bmatrix}\approx\bm{\Psi}=\begin{bmatrix}\bm{\psi}_{1}&\bm{\psi}_{2}&\dots&\bm{\psi}_{N_{\bm{\Omega}}}\end{bmatrix}\approx{\bf{X}}\mathbf{S}\mathbf{V}^{T},$
where $\bm{\Psi}\in\mathbb{R}^{N_{\mathbf{x}}\times N_{\bm{\Omega}}}$, ${\bf{X}}\in\mathbb{R}^{N_{\mathbf{x}}\times r}$, $\mathbf{S}\in\mathbb{R}^{r\times r}$, and $\mathbf{V}\in\mathbb{R}^{N_{\bm{\Omega}}\times r}$. The low-rank factors $\bf{X}$}, $\mathbf{S}$, and $\mathbf{V}$ can be computed using low-rank solvers based on DLR methods [koch2007dynamical] or rank-adaptive approaches [bachmayr2023low, guo2022low]. When the solution exhibits strong low-rank structure, i.e., $r\ll\min(N_{\mathbf{x}},N_{\bm{\Omega}})$, these methods often achieve substantial savings in both memory and computational cost.
In this subsection, we first briefly review existing low-rank iterative solvers for the RTE. We then motivate the present work by highlighting the necessity of, and challenges in, designing highly efficient rank-adaptive low-rank iterative solvers.
2.3.1 Existing low-rank iterative solvers
Low-rank iterative solvers for RTE have been developed to solve linear systems arising from steady-state RTE or implicit time marching of time-dependent RTE. Existing methods can be broadly divided into the following two categories.
Low-rank solvers for the second-order even-parity reformulation of RTE. Methods along this line include [bachmayr2017iterative, guo2025inexact, haut2026efficient]. Rank-adaptive truncated summation has been utilized in low-rank Richardson iteration with an exponential sum preconditioner [bachmayr2017iterative] and in low-rank SI–DSA methods for the second-order RTE [guo2025inexact]. A fixed-rank DLR method for the second-order RTE was introduced in [haut2026efficient].
Sweep-based fixed-rank DLR methods for the first-order formulation of the RTE. An alternative strategy is to directly design a low-rank solver for the first-order RTE based on SI and transport sweeps [peng2023sweep, haut2026efficient]. Peng et al. [peng2023sweep] develop a fixed-rank DLR approach by grouping angular fluxes corresponding to directions within the same octant. This strategy requires modifications to the transport sweeps to solve projected low-rank systems for each angular group. To avoid modifying the transport sweeps, Haut et al. [haut2026efficient] propose a collocation-based DLR method [dektor2025interpolatory] that constructs low-rank solutions via angular sampling using the discrete empirical interpolation method (DEIM) [sorensen2016deim] together with Galerkin projections in the physical space.
2.3.2 Necessity and challenges to design rank-adaptive sweep-based solver
The efficiency and accuracy of low-rank methods are governed by the effective solution rank, which is generally unknown a priori. Therefore, adaptively determining the solution rank is essential for achieving genuine computational savings. Standard rank-adaptive strategies enrich the mode space by expanding its dimension from $r$ to a multiple of $r$, followed by truncation [bachmayr2023low, ceruti2022rank]. For instance, the rank-adaptive DLR method enlarges the space to $2r$ [ceruti2022rank], while the SAT method for the 1D1V Vlasov equation expands it to $3r$. In addition, the truncation step typically involves operations that scale superlinearly in $r$, such as QR or SVD factorizations. When the effective rank required to achieve the desired accuracy or satisfy the convergence criteria of the iterative solver becomes moderately large, such rank-dependent augmentation can introduce substantial computational overhead. In these regimes, the aggressive space augmentation may offset or even negate the potential savings of low-rank methods compared with efficient full-rank solvers. Indeed, for several challenging benchmark problems considered in this work, the effective ranks are relatively high. To reach the SI-DSA stopping criterion $\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{2}\leq 10^{-6}$, the effective rank is approximately $45.7\%$ of the full rank $\min\{N_{\mathbf{x}},N_{\bm{\Omega}}\}$ for the lattice problem (Sec. 4.4), $38\%$ for the variable scattering case (Sec. 4.2), and $43.35\%$ for the pin-cell case (Sec. 4.3). In such cases, even augmenting the mode space from dimension $r$ to $2r$ may render the low-rank solver slower than its full-rank counterpart.
[…]
3 Rank-adaptive sweep-based SI-DSA via mild space augmentation
In this section, we first give an overview of our low-rank SI algorithm in Sec. 3.1, then delve into details of crucial components, including the residual-based greedy random subsampling strategy for angular directions (Sec.3.2), the incremental update of low-rank bases and projected operators (Sec.3.4), and the stopping criteria for the inner loop of the low-rank SI (Sec.3.3). We summarize our contributions and highlight the key differences between the proposed method and existing fixed-rank sweep-based low-rank approaches [peng2023sweep, haut2026efficient] in Sec.3.6. Finally, we outline possible extensions of the method to time-dependent problems in Sec.3.7.
3.1 Algorithm overview of iterative low-rank SI
In this work, we employ a matrix-based low-rank representation for the angular flux:
(13) $\bm{\Psi}=\begin{bmatrix}\bm{\psi}_{1},&\bm{\psi}_{2},&\dots&\bm{\psi}_{N_{\bm{\Omega}}}\end{bmatrix}\approx{\bf{X}}\mathbf{S}\mathbf{V}^{T},$
where ${\bf{X}}\in\mathbb{R}^{N_{\mathbf{x}}\times r}$, $\mathbf{S}\in\mathbb{R}^{r\times r}$ and $\mathbf{V}\in\mathbb{R}^{N_{\bm{\Omega}}\times r}$.
Our low-rank SI seeks an approximate solution of the SI equation (7) in low-rank form, $\bm{\Psi}^{(n)}\approx{\bf{X}}^{(n)}\mathbf{S}^{(n)}(\mathbf{V}^{(n)})^{T}$, via an iterative procedure. In each inner-loop iteration, both the spatial basis and the rank are updated by augmenting the mode space with a small, fixed number of solutions to equation (7) corresponding to selected angular directions. A residual-based greedy random subsampling strategy is then employed to select new angular directions for the subsequent iteration. Upon the convergence of the inner-loop, we construct the low-rank angular basis $\mathbf{V}^{(n)}$ and the coefficient core $\mathbf{S}^{(n)}$. Given ${\bf{X}}^{(n,k-1)}\in\mathbb{R}^{N_{\mathbf{x}}\times r_{k-1}}$, the $k$-th inner-loop iteration obtains ${\bf{X}}^{(n,k)}$ as follows:
  1. 1.
    Transport sweeps: Given $p$ newly sampled angles, $\{\bm{\Omega}_{s^{(k)}_{1}},\dots,\bm{\Omega}_{s^{(k)}_{p}}\}$, we apply transport sweeps to solve
    (14) $({\bf{D}}_{j}+{\mathbf{\Sigma}}_{t})\bm{\psi}_{j}^{(n)}={\mathbf{\Sigma}}_{s}\bm{\phi}^{(n-1)}+\widetilde{\bm{G}}_{j},\quad j=s_{1}^{(k)},\dots,s_{p}^{(k)}.$
    In our numerical experiments, we set $p\leq 3$.
  2. 2.
    Incremental update of low-rank spatial basis and operators. We update ${\bf{X}}^{(n,k)}$ using the modified Gram–Schmidt procedure with selective reorthogonalization (MGS-RO), as described in Alg. 2, to compute its QR factorization incrementally.
    (15) ${\bf{X}}^{(n,k)}\mathbf{R}^{(n,k)}=\begin{bmatrix}{\bf{X}}^{(n,k-1)}&\bm{\psi}^{(n)}_{s^{(k)}_{1}}&\dots&\bm{\psi}^{(n)}_{s^{(k)}_{p}}\end{bmatrix}$,
    where ${\bf{X}}^{(n,k)}\in\mathbb{R}^{N_{\mathbf{x}}\times r_{k}}$ has orthonormal columns, $\mathbf{R}\in\mathbb{R}^{r_{k}\times r_{k}}$ is upper triangular, and the updated rank satisfies $r_{k}=r_{k-1}+p$. If reorthogonalization is not performed, ${\bf{X}}^{(n,k)}$ and ${\bf{X}}^{(n,k-1)}$ satisfy the hierarchical relation ${\bf{X}}^{(n,k)}=\begin{bmatrix}{\bf{X}}^{(n,k-1)}&\mathbf{P}^{(n,k)}\end{bmatrix}$. Leveraging this hierarchical relation, the projection operators used in the sampling step can be updated efficiently in an incremental manner. See Sec. 3.4.
  3. 3.
    Residual-based greedy random subsampling of angular directions. We randomly sample $q>p$ previously unsampled angles to form the candidate set $\mathcal{S}^{(k)}=\{\bm{\Omega}_{t^{(k)}_{1}},\dots,\bm{\Omega}_{t^{(k)}_{q}}\}$. We obtain reduced solutions for angles in $\mathcal{S}^{(k)}$ through Galerkin projections. Based on the corresponding PDE residuals, we greedily select the $p$ angles in $\mathcal{S}^{(k)}$ with the largest residuals $\{\bm{\Omega}_{s^{(k+1)}_{1}},\dots,\bm{\Omega}_{s^{(k+1)}_{p}}\}$. See Sec. 3.2.
  4. 4.
    Inner-loop stopping criteria. If the proposed stopping criteria in Sec. 3.3 are satisfied, we compute $\phi^{(n,*)}$ via numerical integration, and update the low-rank bases and coefficient core through projections and SVD truncation.
Main features of the above low-rank SI are summarized as follows.
  1. 1.
    Sampling in the angular space and Galerkin projection in the physical space. We combine angular sampling with physical-space Galerkin projections to construct low-rank solutions, similar to [haut2026efficient]. This hybrid strategy enables the reuse of existing transport-sweep implementations in a minimally invasive manner; see Sec. 3.6 for details.
  2. 2.
    Mild space augmentation to achieve rank adaptivity. In each inner-loop iteration of our low-rank SI, the mode space is mildly enlarged by adding a small rank-independent number, namely $p$, of snapshots without immediate truncation. A single truncation step is performed only after the inner loop has converged to compress the low-rank representation. As a result, the dimension of the augmented mode space before one-time truncation is only slightly larger than $r$, in contrast to most existing methods that rely on overly aggressive $O(r)$ space augmentation.
  3. 3.
    Delayed construction of $\mathbf{S}^{(n)}$ and $\mathbf{V}^{(n)}$ to enable efficient subsampling. We observe that, with a properly designed stopping criterion, it is unnecessary to explicitly construct $\mathbf{V}^{(n)}$ and $\mathbf{S}^{(n)}$ at every inner-loop iteration of the low-rank SI. This design choice enables an efficient residual-based greedy random subsampling strategy for rank adaptation. The proposed subsampling approach reduces the cost of solving projected systems during angular sampling from $O(N_{\bm{\Omega}}r^{3})$ to $O(qr^{3})$ with $q\ll N_{\bm{\Omega}}$.
  4. 4.
    Incremental update of spatial basis and projected operators. Since no truncation is performed before the convergence of the inner-loop iterations, the adaptively expanded mode space used to construct the low-rank spatial basis possesses a hierarchical structure. Leveraging this structure, we perform efficient incremental updates of both the low-rank spatial basis and the associated Galerkin projected operators, thereby further improving computational efficiency.
With the proposed rank-adaptive low-rank SI, we are now ready to introduce the corresponding low-rank SI-DSA scheme. Each iteration still follows the same two-step procedure as in the full-rank setting: we first apply the low-rank SI to compute a low-rank approximation of the SI equation (7), and then correct the scalar flux to accelerate convergence by solving the DSA equation (8). The outer-loop stopping criterion is still chosen as $\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{2}$ to be sufficiently small. Details of the low-rank SI-DSA are summarized in Alg. 1.
Algorithm 1 Low-rank SI-DSA.
1:Input: Initial scalar flux $\bm{\phi}^{(0)}$, tolerance $\epsilon_{\textrm{Diff}}$, and maximum number of iterations, $N_{\textrm{iter}}$.
2:Initialization: Set $n=1$.
3:while $k\leq N_{\textrm{Iter}}$ do
4:  Apply the low-rank SI following Sec. 3.1 to solve (7) and obtain $\bm{\phi}^{(n,*)}$. Optionally return a rank-$r_{n}$ factorization of the angular flux, ${\bf{X}}^{(n)}\mathbf{S}^{(n)}(\mathbf{V}^{(n)})^{T}$.
5:  if $\|\bm{\phi}^{(n,*)}-\bm{\phi}^{(n-1)}\|_{\infty}\leq\epsilon_{\textrm{SI-SA}}$ then
6:   break
7:  end if
8:  Compute the DSA correction $\delta\bm{\phi}^{(n)}$ by solving a diffusion correction equation, then update the scalar flux as $\bm{\phi}^{(n+1)}=\bm{\phi}^{(n,*)}+\delta\bm{\phi}^{(n)}$.
9:  Set $n\leftarrow n+1$.
10:end while
11:Output: Low-rank angular flux ${\bf{X}}^{(n)}\mathbf{S}^{(n)}(\mathbf{V}^{(n)})^{T}$ and scalar flux $\bm{\phi}^{(n)}$.
3.2 Greedy random subsampling of angles
We present the greedy random subsampling strategy to select angles for rank adaptation in the subsequent iteration. As outlined in Sec. 3.1, instead of considering all angular directions, we first randomly sample $q$ previously unsampeld angles to form the candidate set $\mathcal{S}^{(k)}=\{\bm{\Omega}_{t^{(k)}_{1}},\dots,\bm{\Omega}_{t^{(k)}_{q}}\}$, then compute reduced solutions corresponding to angles in $\mathcal{S}^{(k)}$ through Galerkin projections. Based on the associated residuals, we greedily select $p$ angular directions with the largest residuals for augmentation in the next iteration.
Galerkin projection. Given the updated spatial basis ${\bf{X}}^{(n,k)}$, for an angular direction $\bm{\Omega}_{j}\in\mathcal{S}^{(k)}$ we compute the reduced solution via a Galerkin projection by solving
(16) $\mathbf{A}_{r_{k}}^{(n,k)}\mathbf{c}_{j}^{(n,k)}\triangleq({\bf{X}}^{(n,k)})^{T}({\bf{D}}_{j}+{\mathbf{\Sigma}}_{t}){\bf{X}}^{(n,k)}\mathbf{c}_{j}^{(n,k)}=({\bf{X}}^{(n,k)})^{T}({\mathbf{\Sigma}}_{s}\bm{\phi}^{(n-1)}+\widetilde{\bm{G}}_{j}),$
where $\mathbf{c}_{j}^{(n,k)}\in\mathbb{R}^{r_{k}}$ and $r_{k}=pk$ denotes the rank of the spatial basis ${\bf{X}}^{(n,k)}$. The projected approximation of the angular flux corresponding to these directions is given by $\bm{\psi}_{j}^{(n,k)}\approx{\bf{X}}^{(n,k)}\mathbf{c}_{j}^{(n,k)}$.
Residual-based greedy sampling. After computing the projected solutions, we evaluate the PDE residual of the SI equation (7) for each candidate angular direction in $\mathcal{S}^{(k)}$:
(17) $\bm{r}^{(k)}_{j}={\mathbf{\Sigma}}_{s}\bm{\phi}^{(k)}+\widetilde{\bm{G}}_{j}-({\bf{D}}_{j}+{\mathbf{\Sigma}}_{t}){\bf{X}}^{(n,k)}\mathbf{c}_{j}^{(n,k)},\qquad\Omega_{j}\in\mathcal{S}^{(k)}.$
We then sort the indices of the angles in $\mathcal{S}^{(k)}$ in descending order of their $\ell_{2}$-norms:
$\|\bm{r}^{(k)}_{s^{(k+1)}_{1}}\|_{2}\geq\cdots\geq\|\bm{r}^{(k)}_{s^{(k+1)}_{q}}\|_{2}.$
The first $p$ angular directions corresponding to the largest residuals are selected for augmentation.
Initial sampling. To initialize the inner iteration of the low-rank SI, an initial set of $p$ angular directions $\Omega_{s^{(0)}_{1}},\dots,\Omega_{s^{(0)}_{p}}$ is selected at random. Problem-specific initialization strategies may further enhance efficiency.
Remark 3.1.
Comparison with an alternative DEIM-based sampling strategy. As an alternative to our proposed strategy, one may consider a sampling strategy similar to [dektor2021rank] based on the discrete empirical interpolation method (DEIM) [sorensen2016deim]. Specifically, reduced solutions for all unsampled angular directions can be computed via Galerkin projections by solving (16), yielding the low-rank approximation
(18) $\bm{\Psi}\approx{\bf{X}}^{(n,k)}\begin{bmatrix}\mathbf{c}^{(n,k)}_{1}&\dots&\mathbf{c}^{(n,k)}_{N_{\Omega}}\end{bmatrix}\triangleq{\bf{X}}^{(n,k)}{\bf{C}}^{(n,k)}.$
For previously sampled angles, $\mathbf{c}^{(n,k)}_{j}$ coincides with the corresponding column of $\mathbf{R}^{(n,k)}$ in (15). The matrix ${\bf{C}}^{(n,k)}\in\mathbb{R}^{r_{k}\times N_{\Omega}}$ collects the reduced coefficients of the angular flux in the current spatial basis. An angular basis and a coefficient core can then be obtained via SVD
${\bf{C}}^{(n,k)}=\mathbf{U}^{(n,k)}\mathbf{S}^{(n,k)}(\mathbf{V}^{(n,k)})^{T}.$
New angles are subsequently selected by applying DEIM-induced sampling [sorensen2016deim] to the reduced angular basis $\mathbf{V}^{(n,k)}$.
In challenging benchmark problems, however, we observe that this DEIM-based strategy does not outperform the proposed subsampling approach in terms of sampling quality or the number of required augmentation directions. Moreover, it incurs substantially higher computational cost, since Galerkin projections must be applied for all unsampled angular directions, rather than only for a small candidate set as in the proposed strategy.
3.3 Stopping criteria for the inner-loop of low-rank SI
A commonly used stopping criterion in low-rank iterative solvers (see, e.g., [appelo2025lraa]) measures the difference between two successive low-rank iterates, namely
$\|{\bf{X}}^{(n,k)}\mathbf{S}^{(n,k)}(\mathbf{V}^{(n,k)})^{T}-{\bf{X}}^{(n,k-1)}\mathbf{S}^{(n,k-1)}(\mathbf{V}^{(n,k-1)})^{T}\|_{F}\leq\textrm{threshold},$
which can be computed through low-rank truncated summation without explicitly reconstructing the full matrices. Nevertheless, the procedure [bachmayr2023low, einkemmer2025review] enlarges the low-rank space to dimension $r_{k}+r_{k-1}$ and typically requires a QR or SVD. For the challenging benchmarks considered in Sec. 2.3.2, such augmentation can lead to substantial and unnecessary computational overhead.
To design an effective and efficient stopping criterion, we propose the following strategy, which avoids performing expensive truncated summation at every iteration.
  1. 1.
    When the low-rank SI starts, initialize
    $\bm{\phi}^{(n)}_{\mathrm{old}}:=\bm{\phi}^{(n-1)}.$
  2. 2.
    In the $k$-th inner iteration of the low-rank SI, we first check whether the residuals for all $q$ candidate angles in the greedy random subsampling step are sufficiently small:
    (19) $\max_{\Omega_{j}\in\mathcal{S}^{(k)}}\|\bm{r}_{j}^{(k)}\|_{2}<\epsilon_{\mathrm{res}}.$
  3. 3.
    If (19) holds, we solve the Galerkin projection problem (16) for all previously unsampled angles, yielding the low-rank approximation of the angular flux as in (18). The scalar flux can then be evaluated at cost $O(r_{k}(N_{\mathbf{x}}+N_{\Omega}))$ as
    (20) $\bm{\phi}^{(n,k)}={\bf{X}}^{(n,k)}\bigl({\bf{C}}^{(n,k)}\mathbf{w}\bigr),\qquad\mathbf{w}=(\omega_{1},\dots,\omega_{N_{\Omega}})^{T},$
    where $\mathbf{w}$ collects the normalized quadrature weights satisfying $\sum_{j=1}^{N_{\Omega}}\omega_{j}=1$.
    We then check whether
    (21) $\|\bm{\phi}^{(n,k)}-\bm{\phi}^{(n)}_{\mathrm{old}}\|_{2}\leq\epsilon_{\mathrm{diff}}.$
    If (21) is not satisfied, update
    $\bm{\phi}^{(n)}_{\mathrm{old}}:=\bm{\phi}^{(n,k)}.$
    Otherwise, terminate the inner loop and set
    $\bm{\phi}^{(n,*)}=\bm{\phi}^{(n,k)}.$
Rationale for the stopping criterion. The proposed stopping strategy combines two criteria: the residuals of the SI equation for candidate angles in the sampling step and the change in the scalar flux between successive inner iterations. Note that relying solely on the residuals of candidate angles may lead to premature termination of the inner-loop iterations, which can adversely affect the convergence of the outer SI–DSA iterations. On the other hand, evaluating residuals for all angular directions incurs a prohibitive computational cost of $O(N_{\Omega}N_{\mathbf{x}}r_{k})$. Meanwhile, using only the difference in scalar flux between successive iterations is also insufficiently robust. We numerically observe that, for certain angular directions, the residuals of the SI equation can remain significantly larger than the corresponding change in the scalar flux. In such cases, stopping based solely on the scalar-flux difference may cause premature termination of the inner loop as well, leading to stagnation in the outer-loop convergence. In contrast, the proposed combined criterion achieves both efficiency and robustness across all benchmark tests. The resulting low-rank SI–DSA converges with essentially the same number of outer iterations as its full-rank counterpart.
Construction of $\mathbf{V}^{(n)}$ and $\mathbf{S}^{(n)}$ after convergence. After the inner-loop iterations of the low-rank SI have converged, the angular basis and coefficient core can be obtained via a truncated SVD
${\bf{C}}^{(n,k)}\approx\mathbf{U}^{(n)}\mathbf{S}^{(n)}(\mathbf{V}^{(n)})^{T},$
where $\mathbf{U}^{(n)}\in\mathbb{R}^{r_{k}\times r^{\prime}}$, $\mathbf{S}^{(n)}\in\mathbb{R}^{r^{\prime}\times r^{\prime}}$, $\mathbf{V}^{(n)}\in\mathbb{R}^{N_{\Omega}\times r^{\prime}}$, and $r^{\prime}\leq r_{k}$. The spatial basis is then updated as
${\bf{X}}^{(n)}={\bf{X}}^{(n,k)}\mathbf{U}^{(n)}.$
This construction step is optional, since only the scalar flux $\bm{\phi}^{(n,*)}$ is required to proceed with the outer SI–DSA iterations. Even when the angular flux is needed, for instance, in implicit time marching of time-dependent problems, a single truncation performed after convergence of the outer loop is sufficient.
3.4 Incremental basis and reduced operator update
In the inner-loop iteration of our low-rank SI, the space of spatial modes is progressively enlarged by adding snapshots of angular fluxes for sampled angles. By exploiting this hierarchical structure, the spatial basis ${\bf{X}}^{(n,k)}$ and Galerkin projected operators onto its span can be updated incrementally. Specifically, the orthonormal spatial basis ${\bf{X}}^{(n,k)}$ can be updated through MGS-RO in Alg. 2. Reorthogonalization is selectively applied to mitigate the loss of orthogonality in ${\bf{X}}^{(n,k)}$ caused by numerical error accumulation in the MGS process.
Algorithm 2 Modified Gram–Schmidt with selective reorthogonalization (MGS-RO)
1:Input: Orthonormal spatial basis ${\bf{X}}^{(n,k-1)}$, upper triangular matrix $\mathbf{R}^{(n,k-1)}$, angular snapshots $\bm{\psi}^{(n)}_{s^{(k)}_{1}},\dots,\bm{\psi}^{(n)}_{s^{(k)}_{p}}$, and tolerance $\epsilon_{\mathrm{MGS}}$.
2: Set ${\bf{X}}:={\bf{X}}^{(n,k-1)}$ and $\mathbf{R}:=\mathbf{R}^{(n,k-1)}$.
3:for $l=1:p$ do
4:  Compute $\hat{\mathbf{c}}={\bf{X}}^{T}\bm{\psi}^{(n)}_{s^{(k)}_{l}}$ and $\hat{\bm{\psi}}=\bm{\psi}^{(n)}_{s^{(k)}_{l}}-{\bf{X}}\hat{\mathbf{c}}$.
5:  Define $\tilde{\mathbf{c}}=\bigl(\hat{\mathbf{c}}^{T},\ \|\hat{\bm{\psi}}\|_{2}\bigr)^{T}$.
6:  Update ${\bf{X}}:=\begin{bmatrix}{\bf{X}}&\hat{\bm{\psi}}/\|\hat{\bm{\psi}}\|_{2}\end{bmatrix}$ and $\mathbf{R}:=\begin{bmatrix}\mathbf{R}&\tilde{\mathbf{c}}\end{bmatrix}$.
7:end for
8:if $\bigl|{\bf{X}}(:,1)^{T}{\bf{X}}(:,\mathrm{end})\bigr|>\epsilon_{\mathrm{MGS}}$ then
9:  Recompute a QR factorization:
${\bf{X}}^{(n,k)}\mathbf{R}^{(n,k)}=\mathrm{qr}\!\left(\begin{bmatrix}{\bf{X}}^{(n,k-1)}&\bm{\psi}^{(n)}_{s^{(k)}_{1}}&\dots&\bm{\psi}^{(n)}_{s^{(k)}_{p}}\end{bmatrix}\right)$.
10:else
11:  ${\bf{X}}^{(n,k)}:={\bf{X}}$ and $\mathbf{R}^{(n,k)}:=\mathbf{R}$.
12:end if
If reorthogonalization is not performed, $\small{{\bf{X}}^{(n,k)}=\begin{bmatrix}{\bf{X}}^{(n,k-1)}&\mathbf{P}^{(n,k)}\end{bmatrix}}$} with $\mathbf{P}^{(n,k)}\in\mathbb{R}^{N_{\mathbf{x}}\times p}$. Then, the projected discrete total cross section can be incrementally updated as (dropping the superscript index (n,k1)(n,k-1) for brevity):
(22) $({\bf{X}}^{(n,k)})^{T}{\mathbf{\Sigma}}_{t}{\bf{X}}^{(n,k)}=(\begin{bmatrix}{\bf{X}}&\mathbf{P}\end{bmatrix})^{T}{\mathbf{\Sigma}}_{t}\begin{bmatrix}{\bf{X}}&\mathbf{P}\end{bmatrix}=\begin{bmatrix}\underbrace{{\bf{X}}^{T}{\mathbf{\Sigma}}_{t}{\bf{X}}}_{\textrm{Known}}&{\bf{X}}^{T}({\mathbf{\Sigma}}_{t}\mathbf{P})\\ (\mathbf{P}^{T}{\mathbf{\Sigma}}_{t}){\bf{X}}&\mathbf{P}^{T}{\mathbf{\Sigma}}_{t}\mathbf{P}\end{bmatrix}$.
The projections of other operators can be updated in the same manner. In particular, the projection of the transport operator for the angle $\bm{\Omega}_{j}$, given by $({\bf{X}}^{(n,k)})^{T}{\bf{D}}_{j}{\bf{X}}^{(n,k)}$, can be calculated as a linear combination of the projections of ${\bf{D}}_{x}^{\pm}$ and ${\bf{D}}_{y}^{\pm}$ defined in (6b).
3.5 Computational cost of each inner-loop iteration in low-rank SI
Here, we summarize the computational cost of the $k$-th inner-loop iteration in low-rank SI.
  1. 1.
    Transport sweeps: Solving the SI equation (7) for the $p$ newly sampled directions requires $O(pN_{\mathbf{x}})$ operations.
  2. 2.
    Basis and operators update: Without reorthogonalization, the incremental update of the spatial basis and projected operators incurs a cost of $O(N_{\mathbf{x}}r_{k-1}p)$. If reorthogonalization is performed, the cost scales as $O(N_{\mathbf{x}}r_{k}^{2})$.
  3. 3.
    Greedy random subsampling: Solving the projected systems for the $q$ candidate angles and computing their corresponding residuals incur a cost of $O(qr_{k}^{3})$ and $O(N_{\mathbf{x}}r_{k}q)$, respectively.
  4. 4.
    Convergence monitoring: The cost of evaluating stopping criteria depends on the residual magnitude. If the maximum residual among candidate angles exceeds the tolerance, no further cost is incurred. Otherwise, generating the coefficient matrix $\mathbf{C}^{(n,k)}$ requires additional $O((N_{\mathbf{\Omega}}-q-r_{k})r_{k}^{3})$ operations, and updating the scalar flux costs $O((N_{\mathbf{x}}+N_{\mathbf{v}})r_{k})$.
We emphasize that, as observed in our numerical experiments, projections for the full set of unsampled angular directions, together with the subsequent angular integration, are performed only in the final few iterations before convergence, thereby significantly reducing the average per-iteration cost. Moreover, reorthogonalization is not required in most inner iterations.
3.6 Contributions and relation to prior work
Fixed-rank sweeping-based DLR methods for the first-order formulation of the RTE have been developed in [peng2023sweep, haut2026efficient]. To the best of our knowledge, the present work introduces the first rank-adaptive sweeping-based low-rank SI framework for the RTE. In contrast to classical rank-adaptation strategies that rely on $O(r)$ space augmentation, we propose a novel and efficient rank-adaptive methodology that achieves mild basis enrichment through residual-based greedy random subsampling of angular directions that only adds a constant number of basis per inner-loop iteration.
Although both our method and the concurrent $S_{N}$-like collocation-based DLR method in [haut2026efficient] construct low-rank approximations via angular sampling and physical-space projections, the underlying algorithmic structures differ fundamentally. The method in [haut2026efficient] is designed for implicit time marching of time-dependent thermal radiation transfer problems. At each time step, a fixed number of angular directions is selected using a DEIM-based sampling procedure applied to the low-rank angular basis from the previous time step, followed by the application of SI–DSA to a modified $S_{N}$ system determined by those angles. Thus, angular sampling is performed outside the SI–DSA iteration in [haut2026efficient]. In contrast, our method integrates adaptive angular sampling directly into each inner SI iteration. This integration is essential for enabling rank adaptivity with only mild space augmentation.
Compared with both our approach and [haut2026efficient], the method in [peng2023sweep] performs projections in the angular space. Let $s$ denote the number of spatial DOFs per element. The sweeping procedure in [peng2023sweep] requires solving a modified $rs\times rs$ system in each element, whereas both [haut2026efficient] and our approach retain the original $s\times s$ system arising directly from the DG discretization. Consequently, existing transport sweep implementations can be reused without modification in [haut2026efficient] and in our framework. Meanwhile, [haut2026efficient] and our framework differ in how existing SI–DSA components are incorporated. The method in [haut2026efficient] retains the entire SI–DSA framework. In contrast, our framework preserves the classical transport sweep and DSA preconditioner, but replaces the full-rank SI iteration with the proposed low-rank SI.
To further place our work in the broader context of efficient rank adaptation, we note that aggressive basis augmentation was also mitigated in [hu2022adaptive] and later integrated into spatial domain decomposition in [brunner2026domain]. In that approach, a small number of random orthogonal basis vectors, determined by the rank of the boundary conditions, are added at the beginning of each time step. In contrast, our method targets iterative solvers and enriches the basis through error-indicated sampling rather than totally random padding.
3.7 Future extensions to implicit time marching
The proposed low-rank SI-DSA can be straightforwardly extended to solve linear systems arising from implicit time marching, as they can be viewed as modified steady-state problems in each time step. For example, each time step of the backward Euler scheme can be interpreted as solving a steady-state problem with a modified discrete total cross section $\widetilde{{\mathbf{\Sigma}}}_{t}={\mathbf{\Sigma}}_{t}+\frac{1}{\Delta t}{\bf{M}}$, where $\bf{M}$} denotes a mass matrix and $\Delta t$ is the time step size. As the primary focus of this work is the development of efficient iterative solvers, a detailed investigation of such time-dependent extensions is deferred to future work.
4 Numerical examples
We demonstrate the performance of our method via a series of benchmark 2D2V tests. Throughout this section “LR” and “FR” refer to low-rank and full-rank solutions, respectively.
In all tests, we use piecewise linear polynomial basis for the DG discretization. A fully-consistent DSA preconditioner is employed [Adams2002FastIM] (see Appx. B of [peng2024reduced] for details). The diffusion correction equation in DSA is solved using the conjugate gradient method preconditioned by a V-cycle algebraic multigrid method with the relative residual tolerance set to $10^{-12}$. For comparison, the stopping criteria is chosen as $\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{\infty}\leq\epsilon_{\textrm{SI-SA}}=10^{-6}$ for both the full-rank SI-DSA and the outer-loop iteration of low-rank SI-DSA (see Alg. 1 for details). In addition, a zero initial guess is used for both methods. The initial angle samples for the first inner-loop iteration of low-rank SI-DSA are chosen randomly.
In our low-rank SI-DSA, unless otherwise specified, we set $\epsilon_{\textrm{res.}}=\epsilon_{\textrm{diff.}}=10^{-7}$ in the stopping criteria for the inner-loop iterations of low-rank SI, and choose $p=1,q=8$ in the residual-based greedy random subsampling of angles. In other words, by default, we select one angle sample from $8$ random candidates. In addition, we construct the low-rank core and the angular flux by truncating at the rank $r$ satisfies $r=\arg\min_{l}\sum_{i=1}^{l}s_{i}/\sum_{i=1}^{r_{n}}s_{i}\leq\epsilon_{\textrm{SVD}}$, where $s_{i}$ are singular values in the SVD of $C^{(n)}$ after the inner-loop convergence and $\epsilon_{\textrm{SVD}}=10^{-8}$.
To quantify the performance of our method, we define metrics including the speedup, compression ratio and inner-loop oversampling ratio as follows.
[…]
(25) inner-loop oversampling ratio $=\frac{\textrm{Number of Sampled Angles}-\textrm{Solution Rank}}{\textrm{Solution Rank}}.$
Before delving into details, we outline numerical tests performed and their purposes. In Sec. 4.1, for a homogeneous medium ranging from weak to strong scattering, we perform a refinement study to investigate the influence of mesh resolution and the effectiveness of our method in different regimes. In Sec. 4.2, we simulate a multiscale variable scattering problem and investigate the influence of $p$ and $q$ in our sampling step. In Sec. 4.3, a pin-cell problem with strong material discontinuities is solved repeatedly to demonstrate the robustness of the method with respect to randomness in the subsampling step. In Sec. 4.4, a lattice problem is solved using both the proposed method and a modified version to illustrate the necessity of the incremental update of the spatial basis and projection operators.
4.1 Homogeneous medium in different regime
We consider the computational domain $1,1]^{2}$ with no absorption, $\sigma_{s}(x,y)=0.1,1,10,100$ and zero inflow boundary conditions. An isotropic Gaussian source $G(x,y)=\exp(-100(x^{2}+y^{2}))$ is imposed. We perform a refinement study using a $16L\times 16L$ rectangular mesh in the physical space and the CL($8L,4L$) quadrature in the angular space for $L=2,3,4,5$. For comparison, the reference scalar flux obtained by the full-rank SI-DSA with the resolution corresponding to $L=5$ is presented in Fig. 1 on a logarithmic scale.
The main goal of this example is to investigate the influence of $\sigma_{s}$ and the mesh resolution on our low-rank SI-DSA. The results are summarized in Tab. 1. Main observations are as follows.
Accuracy and number of iterations for convergence. Regardless of the mesh resolution and the strength of the scattering effect, both the low-rank SI-DSA and full-rank SI-DSA converge within exactly the same number of iterations. Moreover, the difference between the scalar fluxes computed by low-rank and full-rank SI-DSA is at most on the order of $10^{-7}$, which is at least one order of magnitude smaller than $\epsilon_{\textrm{SI-SA}}$ in the outer-loop stopping criterion of SI-DSA. These observations demonstrate the correctness and effectiveness of our low-rank SI-DSA.
Influence of mesh resolution. For all $\sigma_{s}$, increasing the mesh resolution leads to greater memory savings, and the speedup of the low-rank SI-DSA over the full-rank counterpart generally increases as well.
Influence of scattering strength $\sigma_{s}$.
  1. 1.
    Convergence. When $\sigma_{s}=0.1$ (transport regime) or $100$ (diffusion regime), both full-rank and low-rank SI-DSA converges within 454-5 iterations. In contrast, more iterations are required for convergence in the intermediate regime ($\sigma_{s}=1,10$).
  2. 2.
    Rank. The solution rank almost grows linearly with respect to the grid resolution in the transport regime ($\sigma_{s}=0.1$), scales roughly as $O(L^{1.5})$ in the intermediate regime ($\sigma_{s}=1,10$), and almost remains constant once the resolution is sufficiently fine in the diffusion regime ($\sigma_{s}=100$). The resolution dependence of the rank in the transport-dominated and intermediate regimes is consistent with observations in our previous work [guo2025inexact].
Computational savings compared to full-rank solver. The low-rank SI-DSA consistently achieves computational time and memory savings over the full-rank SI-DSA across all configurations in this test. The largest savings occur in the diffusion regime ($\sigma_{s}=100$). In particular, when $\sigma_{s}=100$ and $L=5$, our low-rank SI-DSA gains more than $12\times$ reductions in both the computational time and memory.
$L$ Rank $\|\bm{\psi}_{\textrm{LR}}-\bm{\psi}_{\textrm{FR}}\|_{2}$ $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ C-R FR-Iter LR-Iter Speedup
22 $56$ $3.62\times 10^{-8}$ $4.28\times 10^{-9}$ $45.13\%$ 55 55 $1.70\times$
33 $93$ $1.38\times 10^{-7}$ $1.07\times 10^{-8}$ $33.30\%$ 55 55 $2.77\times$
44 $127$ $2.31\times 10^{-7}$ $7.99\times 10^{-9}$ $25.58\%$ 55 55 $3.24\times$
55 $160$ $6.60\times 10^{-7}$ $1.41\times 10^{-8}$ $20.63\%$ 55 55 $4.86\times$
(a) $\sigma_{s}=0.1$.
$L$ Rank $\|\bm{\psi}_{\textrm{LR}}-\bm{\psi}_{\textrm{FR}}\|_{2}$ $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ C-R FR-Iter LR-Iter Speedup
22 $64$ $2.31\times 10^{-9}$ $1.22\times 10^{-9}$ $51.57\%$ 88 88 $1.37\times$
33 $123$ $2.84\times 10^{-7}$ $1.84\times 10^{-8}$ $44.05\%$ 88 88 $1.86\times$
44 $175$ $1.08\times 10^{-7}$ $7.31\times 10^{-9}$ $35.25\%$ 88 88 $2.29\times$
55 $234$ $6.70\times 10^{-7}$ $1.57\times 10^{-8}$ $30.17\%$ 88 88 $2.44\times$
(b) $\sigma_{s}=1$
$L$ Rank $\|\bm{\psi}_{\textrm{LR}}-\bm{\psi}_{\textrm{FR}}\|_{2}$ $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ C-R FR-Iter LR-Iter Speedup
22 $64$ $3.23\times 10^{-7}$ $2.96\times 10^{-7}$ $51.57\%$ 88 88 $1.55\times$
33 $141$ $1.28\times 10^{-7}$ $5.49\times 10^{-8}$ $50.49\%$ 88 88 $1.04\times$
44 $194$ $6.59\times 10^{-8}$ $6.07\times 10^{-9}$ $39.08\%$ 88 88 $1.57\times$
55 $229$ $7.05\times 10^{-8}$ $4.37\times 10^{-9}$ $29.52\%$ 88 88 $1.64\times$
(c) $\sigma_{s}=10$
$L$ Rank $\|\bm{\psi}_{\textrm{LR}}-\bm{\psi}_{\textrm{FR}}\|_{2}$ $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ C-R FR-Iter LR-Iter Speedup
22 $53$ $7.11\times 10^{-8}$ $7.06\times 10^{-8}$ $2.71\% 44$ 44 $1.76\times$
33 $61$ $1.80\times 10^{-7}$ $1.79\times 10^{-7}$ $21.84\%$ 44 44 $3.73\times$
44 $63$ $1.71\times 10^{-7}$ $1.70\times 10^{-7}$ $12.69\%$ 44 44 $6.71\times$
55 $63$ $1.93\times 10^{-7}$ $1.91\times 10^{-7}$ $8.12\%$ 55 55 $13.88\times$
(d) $\sigma_{s}=100$
Table 1: Results for the homogeneous medium problem in Sec. 4.1. C-R: compression ratio. FR/LR-Iter: number of source iterations required by convergence for the full/low-rank SI-DSA. $L$: refinement level corresponding to $(N_{x},N_{y},N_{\theta},N_{\Omega_{z}})=(16L,16L,8L,4L)$.
Figure 1: Full-rank reference scalar flux for the homogeneous problem in Sec. 4.1 obtained with $(N_{x},N_{y},N_{\theta},N_{\bm{\Omega}_{z}})=(80,80,40,20)$. Top left: $\sigma_{s}=0.1$. Top right: $\sigma_{s}=1$. Bottom left: $\sigma_{s}=10$. Bottom right: $\sigma_{s}=100$.
4.2 Multiscale variable scattering problem
We simulate a multiscale, variable scattering problem on the computational domain $1,1]^{2}$ with vacuum boundary conditions, no absorption, and a Gaussian source $G(x,y)=\exp(-100(x^{2}+y^{2}))$. The scattering cross section (illustrated in Fig. 2) is defined as
(26) $\sigma_{s}(x,y)=\begin{cases}99.9r^{4}(r^{2}-2)^{2}+0.1;,\quad\text{if}\;r=\sqrt{x^{2}+y^{2}}\leq 1,\\ 100,\quad\text{otherwise}.\end{cases}$
This scattering cross section increases smoothly from $0.1$ in the center to $100$ towards the boundary, implying a smooth transition from the transport regime in the center to the scattering regime near the boundary. Hence, this example exhibits significant multiscale effects. We partition physical domain with a $80\times 80$ mesh and employ the CL($40,20$) quadrature in angular space. The scalar fluxes obtained by full-rank and low-rank SI-DSA are presented on a logarithmic scale in Fig. 2. No visual difference is observed.
We further test the robustness of our low-rank SI-DSA with respect to the hyper-parameters in the angular sampling step, namely $p$ (the number of angles selected each iteration) and $q$ (the number of randomly sampled candidate angles). The results are summarized in Tab. 2(b). Our main observations are as follows.
Robust accuracy and convergence against various $p$ and $q$. Across all choices of $p$ and $q$, our low-rank SI-DSA converges in the same number of iterations as the full-rank SI-DSA. Moreover, the difference between the scalar fluxes produced by the low-rank and full-rank methods, $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ is on the order of $10^{-7}$-$10^{-8}$, at least one order of magnitude smaller than the threshold $\epsilon_{\textrm{SI-SA}}=10^{-6}$ used in the stopping criterion of SI-DSA, i.e. $\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{\infty}\leq\epsilon_{\textrm{SI-SA}}$ .
Robustness against $p$ and $q$. We observe that, for various $p$ and $q$, the low-rank SI-DSA consistently reduces the DOFs required to store the angular flux by approximately $38\%$. Moreover, the reduction in computational time is around $1.20\times$ for most cases, except for $1.05\times$ with $(p,q)=(1,16)$. The slightly smaller speedup in this case may be attributed to the increased computational cost in the sampling step.
Mild space augmentation. In Fig. 3, we present the effective rank and the oversampling ratio during low-rank SI for various $p$ and $q$. With $p=1$, the oversampling ratio remains below $5\%$ for all choices of $q$. For fixed $q=8$, increasing the number of sampled angles per iteration leads to a higher oversampling ratio, but it never exceeds $20\%$. These observations verifies the effectiveness of the proposed incremental strategy in preventing overly aggressive space augmentation during rank adaptation. We emphasize that a rank adaptive method using the standard $O(r)$ space augmentation may even become slower than the full-rank solver, since the solution rank is close to $38\%$ of the full rank under the present setup.
Comparison between the rank of low-rank solutions and the effective rank of the full-rank solution. We compute the effective rank of the full-rank solution by rewriting it in the matrix format as in (12) and performing truncated SVD with the same truncation tolerance as in the low-rank method, i.e., $\epsilon_{\textrm{SVD}}=10^{-8}$. The resulting effective rank is 318, which is close to the ranks of the low-rank solutions ranging from 281 to 315.
Figure 2: Configuration of scattering cross section and scalar fluxes obtained by full-rank and low-rank SI-DSA for the variable scattering example in Sec. 4.2. Left: $\sigma_{s}(x,y)$. Middle: $\phi$ obtained by full-rank SI-DSA. Right: $\phi$ obtained by low-rank SI-DSA.
$q$ Rank $\|\bm{\psi}_{\textrm{LR}}-\bm{\psi}_{\textrm{FR}}\|_{2}$ $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ C-R FR-Iter LR-Iter Speedup
44 $281$ $3.90\times 10^{-6}$ $1.06\times 10^{-7}$ $36.22\%$ 77 77 $1.17\times$
88 $298$ $6.29\times 10^{-7}$ $8.87\times 10^{-8}$ $38.42\%$ 77 77 $1.19\times$
$12$ $297$ $29\times 10^{-7} 1.1$$14\times 10^{-7}$ $38.29\%$ 77 77 $1.22\times$
$16$ $299$ $5.49\times 10^{-8}$ $3.58\times 10^{-8}$ $38.54\%$ 77 77 $1.05\times$
(a) Results for $p=1$ and different $q$.
$p$ Rank $\|\bm{\psi}_{\textrm{LR}}-\bm{\psi}_{\textrm{FR}}\|_{2}$ $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ C-R FR-Iter LR-Iter Speedup
11 $298$ $6.29\times 10^{-7}$ $8.87\times 10^{-8}$ $38.42\%$ 77 77 $1.19\times$
22 $298$ $3.53\times 10^{-6}$ $1.08\times 10^{-7}$ $38.42\%$ 77 77 $1.29\times$
33 $315$ $2.02\times 10^{-6}$ $2.89\times 10^{-7}$ $40.61\%$ 77 77 $1.18\times$
(b) Results for different $p$ and fixed $q=8$
Table 2: Results for the variable scattering example in Sec. 4.2. $p$: angle sampled added each inner-loop iteration. $q$: number of randomly drawn candidate samples angles. C-R: compression ratio. FR/LR-Iter: number of iterations for convergence required by full/low-rank SI-DSA.
Figure 3: Oversampling ratio during low-rank SI for the variable scattering problem. Left: fixed $p=1$ and different $q$. Right: fixed $q=8$ and different pp.
4.3 Pin-cell problem
We consider a pin-cell problem with strong material discontinuities and multiscale effects, a classical benchmark for testing iterative solvers and preconditioners. The computational domain is $1,1]^{2}$. There is no absorption. A Gaussian source, $G(x,y)=\exp(-100(x^{2}+y^{2}))$, and zero inflow boundary conditions are imposed. The scattering cross section (configuration illustrated in Fig. 4) is defined as
(27) $\sigma_{s}(x,y)=\begin{cases}0.1,\;\text{if}\;|x|\leq 0.5\;\text{and}\;|y|\leq 0.5,\\ 100,\;\text{otherwise}.\end{cases}$
The center of the domain lies in the transport-dominated regime ($\sigma_{s}=0.1$), while the outer region is in the diffusion-dominated regime ($\sigma_{s}=100$), resulting in a 1000-fold difference in scattering strength. The domain is discretized with an $80\times 80$ mesh in physical space and the CL(32,16) quadrature in angular space. Fig. 4 shows the scattering configuration and the scalar fluxes computed by the low-rank and full-rank SI-DSA for comparison.
The inner loop of our low-rank SI involves random subsampling of angles, with the initial angles randomly sampled in the first iteration. To assess robustness with respect to randomness, we solve the problem eight times using different random seeds determined by system time. The results are summarized in Tab. 3, and the main observations are as follows.
Effectiveness and robustness with respect to the randomness in the algorithm. From the accuracy perspective, the difference between the solutions produced by the low-rank SI-DSA and the full-rank SI-DSA is consistently much smaller than $\epsilon_{\textrm{SI-SA}}=10^{-6}$. In $7$ out of $8$ runs, our low-rank SI-DSA converges with exactly the same number of iterations as its full-rank counterpart, while one run requires one more iteration.
The solution ranks for all runs of the low-rank SI-DSA are close to $42\%$ of the full rank. Moreover, even in this challenging scenario, thanks to the efficiency of the proposed sampling strategy and the mild space augmentation in the rank-adaptation step, the method achieves at least a $1.11\times$ speedup and an average speedup of $1.17\times$. In the best run, a speedup of approximately $1.40\times$ is obtained.
Figure 5 shows the convergence history and the oversampling ratio during the low-rank SI iterations for the last run. Except for the second-to-last iteration, the change of the scalar flux in the $k$-th iteration, $\|\bm{\phi}^{(k)}-\bm{\phi}^{(k-1)}\|_{2}$, for the low-rank and full-rank SI-DSA almost coincide. The oversampling ratio remains below 9% throughout all iterations. The observations for the other runs are similar and are therefore omitted.
Finally, we compute the effective rank of the full-rank angular flux by applying truncated SVD to its matrix representation with the same truncation tolerance as in the low-rank method. The effective rank of the full-rank solution is 222, while the ranks of the low-rank solutions range from 207 to 218, with an average of 212.63.
Figure 4: Configuration of scattering cross section and scalar fluxes obtained by full-rank and low-rank SI-DSA for the pin-cell example in Sec. 4.2. Left: configuration of $\sigma_{s}(x,y)$, white region corresponding to $\sigma_{s}=0.1$ and black region corresponding to $\sigma_{s}=100$. Middle: $\phi$ obtained by full-rank SI-DSA. Right: $\phi$ obtained by low-rank SI-DSA.
Figure 5: Convergence history, effective rank and oversampling ratio during the low-rank SI for the pin-cell problem in Sec. 4.3. Left: convergence history. Right: oversampling ratio.
[…]
Solution rank $212.63$ ($\pm 3.00$) $207$ $218$
$\|\psi_{\text{LR}}-\psi_{\text{FR}}\|_{2}$ $2.08(\pm 1.73)\times 10^{-7}$ $9.47\times 10^{-8}$ $6.53\times 10^{-7}$
$\|\phi_{\text{LR}}-\phi_{\text{FR}}\|_{2}$ $1.14(\pm 1.49)\times 10^{-7}$ $2.29\times 10^{-8}$ $5.05\times 10^{-7}$
[…]
FR-Iter $21.00$ ($\pm 0.00$) 2121 2121
LR-Iter $21.13$ ($\pm 0.33$) 2121 $22$
Table 3: Statistics for the low-rank SI-DSA solving the pin-cell problem in Sec. 4.3 computed from the results of $8$ random runs. Mean: mean value. SD: standard deviation. FR/LR-Iter: number of iterations required by the convergence of the full/low-rank SI-DSA.
4.4 Lattice problem
Our final test is a lattice problem with zero inflow boundary conditions, and its configuration is shown in Fig. 6. The black region corresponds to pure absorption with $(\sigma_{a},\sigma_{s})=(100,0)$, while the white and orange regions correspond to pure scattering with $(\sigma_{a},\sigma_{s})=(0,1)$. A constant source $G(x,y)=1$ is imposed in the orange region. The computational domain $0,5]^{2}$ is discretized using an $80\times 80$ mesh in physical space and the CL(32,16) quadrature in angular space. The scalar fluxes obtained by the low-rank and full-rank SI-DSA are shown on a logarithmic scale in Fig. 6 and match each other well.
This example is used to demonstrate the necessity of the incremental update of the spatial basis and the projected reduced operators. In particular, we compare the performance of our method with incremental updates against a modified version that replaces MGS-RO with direct QR decomposition in the basis update. We note that the incremental update of projection operators relies on the hierarchical structure of the spatial basis; therefore, incremental updates of the projected operators cannot be applied after this replacement.
Necessity of incremental update. As shown in the left plot of Fig. 7, the histories of $\|\bm{\phi}^{(k)}-\bm{\phi}^{(k-1)}\|_{2}$ for the low-rank and full-rank SI-DSA almost coincide. As reported in Tab. 4, the proposed low-rank SI-DSA achieves approximately $45.22\%$ compression in DOFs required to store the angular flux, converges in the same number of iterations as the full-rank SI-DSA, and achieves a $1.15\times$ speedup over it.
Although replacing the incremental update with direct QR decomposition does not affect the iteration count, the leading digits of the error compared with the full-rank solution and the memory savings remain essentially unchanged. However, the resulting low-rank solver becomes significantly slower, in fact approximately $1.53\times$ slower than the full-rank method.
Verification of mild augmentation. As shown in the right panel of Fig. 6, the iterative incremental update of the spatial basis effectively avoids aggressive augmentation during the low-rank SI iterations, resulting in less than $11.7\%$ over-augmentation in almost all iterations, with only one iteration exhibiting an oversampling ratio of $15.5\%$.
Figure 6: Configuration of scattering cross section and scalar fluxes obtained by full-rank and low-rank SI-DSA for the Lattice example in Sec. 4.2. We note that negative scalar flux is generated by upwind DG method in this example, and we round negative values to $10^{-16}$ for plotting purposes. Left: configuration of $\sigma_{s}(x,y)$, white and orange region pure scattering, black region pure absorption and a constant source imposed in the orange region. Middle: $\phi$ obtained by full-rank SI-DSA. Right: $\phi$ obtained by low-rank SI-DSA.
Figure 7: Convergence history, effective rank and oversampling ratio during low-rank source iterations for the lattice problem in Sec. $4.$ Left: convergence history. Right: oversampling ratio.
Rank $\|\bm{\psi}_{\textrm{LR}}-\bm{\psi}_{\textrm{FR}}\|_{2}$ $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ C-R FR-Iter LR-Iter Speedup
MGS $227$ $4.63\times 10^{-7}$ $3.27\times 10^{-8}$ $45.22\%$ $11$ $11$ $1.15\times$
QR $227$ $4.63\times 10^{-7}$ $3.27\times 10^{-8}$ $45.22\%$ $11$ $11$ $0.65\times$
Table 4: Results for the lattice problem in Sec. $4.$ MGS: modified Gram-Schimdt with selective reorthogonalziation to update spatial basis and incremental update of projected operators. QR: using direct QR decomposition to update spatial basis. FR/LR-Iter: number of iterations required for convergence by the full/low-rank SI-DSA.
5 Conclusions
We develop an efficient, rank-adaptive, sweeping-based low-rank SI-DSA for solving the RTE. Beyond introducing rank adaptivity into sweeping-based iterative solvers for the RTE, the main contribution lies in the design of a novel and efficient rank-adaptation strategy that updates the basis with only mild space augmentation. Even for challenging problems with stringent outer-loop convergence criteria, i.e.,
$\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{\infty}\leq 10^{-6},$
whose solution ranks are approximately $35\%–45\%$ of the full rank, the proposed method produces highly accurate low-rank solutions. The differences from the full-rank solutions are only $O(10^{-7})–O(10^{-8})$, while achieving savings in both memory usage and computational time. When the solution is intrinsically low-rank, e.g., in the diffusion-regime test in Sec. 4.1, the method achieves more than $12\times$ speedup and memory savings. Moreover, the proposed method is shown to be robust with respect to hyperparameters and the inherent randomness in the subsampling step.
In future work, we plan to extend the proposed method to more complex problems involving multiple energy groups, thermal radiation, anisotropic scattering, and time dependence. We also aim to integrate our method with more efficient or robust preconditioners, such as the quasi-diffusion method in [olivier2023family] and data-driven SA preconditioners [mcclarren2022data, peng2024reduced, tang2025synthetic].
Acknowledgments
W. Guo was partially supported by the Air Force Office of Scientific Research FA9550-18-1-0257. Z. Peng was partially supported by the Hong Kong Research Grants Council grants Early Career Scheme 26302724 and General Research Fund 16306825.
Appendix A Chebyshev-Legendre quadrature
The normalized CL quadrature rule, denoted by $\mathrm{CL}(N_{\theta},N_{\bm{\Omega}_{z}})$, is the tensor product of the normalized $N_{\bm{\Omega}_{z}}$-point Gauss–Legendre quadrature on [-1,1], $\{(\bm{\Omega}_{z,j},\omega_{z,j})\}_{j=1}^{N_{\bm{\Omega}_{z}}}$ with $\sum_{j=1}^{N_{\bm{\Omega}_{z}}}\omega_{j}=1$, and the normalized $N_{\theta}$-point Chebyshev quadrature on the unit circle
(28) $\left\{(\theta_{j},\omega_{\theta,j}):\;\theta_{j}=\frac{2j\pi}{N_{\theta}}-\frac{\pi}{N_{\theta}}\,\text{and}\,\omega^{\phi}_{j}=\frac{1}{N_{\theta}},\;j=1,\dots,N_{\theta}\right\}.$
Taking the tensor product of these two quadrature rules, we obtain the quadrature points and the corresponding quadrature weights of the CL($N_{\theta}$$N_{\bm{\Omega}_{z}}$) rule:
(29a) $\bm{\Omega}_{j}=\left(\bm{\Omega}_{j,x},\bm{\Omega}_{j,y},\bm{\Omega}_{j,z}\right)=\left(\cos(\theta_{j_{1}})\sqrt{1-\bm{\Omega}_{z,j_{2}}^{2}},\sin(\theta_{j_{1}})\sqrt{1-\bm{\Omega}_{z,j_{2}}^{2}},\bm{\Omega}_{z,j_{2}}\right),$
(29b) $\omega_{j}=\omega_{\theta,j_{1}}\omega_{\bm{\Omega}_{z},j_{2}},$
where $j=(j_{1},j_{2})$, $j_{1}=1,\ldots,N_{\theta}$, $j_{2}=1,\ldots,N_{\bm{\Omega}_{z}}$. Note that if $N_{\theta}=2N$ and $N_{\bm{\Omega}_{z}}=N$, then the quadrature rule $\mathrm{CL}(N_{\theta},N_{\bm{\Omega}_{z}})$ exactly integrates normalized polynomials of degree up to $2N$ on the unit sphere. Moreover, CL quadrature rule is symmetric: if $\bm{\Omega}_{j}$ is a quadrature point, so is $\bm{\Omega}_{j}$.
Appendix B Definitions in the matrix-vector formulation of $S_{N}$ upwind DG scheme
Recall that $\{\eta_{i}\}_{i=1}^{N_{\mathbf{x}}}$ forms an orthonormal basis of $U_{h}^{K}$. For each basis $\eta_{i}$, we also assume that it is compactly supported in one spatial element $\mathcal{T}_{a_{i},b_{i}}$. This definition implies that $\eta_{i}(x_{\frac{1}{2}}^{-},y)=\eta_{i}(x_{N+\frac{1}{2}}^{+},y)=\eta_{i}(x,y_{\frac{1}{2}}^{-})=\eta_{i}(x,y_{N+\frac{1}{2}}^{+})=0$. Following the Matlab notation, the $(k,l)$-th elements of ${\bf{D}}_{x}^{\pm}$, ${\mathbf{\Sigma}}_{t}$, ${\mathbf{\Sigma}}_{s}$ in equation (6) are defined as
(30a) ${\bf{D}}^{\pm}_{x}(k,l)$ $=\sum_{a=1}^{N_{x}}\sum_{b=1}^{N_{y}}\Big(-\int_{\mathcal{T}_{a,b}}\partial_{x}\eta_{k}\eta_{l}d\mathbf{x}-\int_{y_{b-\frac{1}{2}}}^{y_{b+\frac{1}{2}}}\left(\eta_{k}(x_{a+\frac{1}{2}}^{+},y)-\eta_{k}(x_{a+\frac{1}{2}}^{-},y)\right)\eta_{l}(x_{a+\frac{1}{2}}^{\pm},y)dy\Big),$
(30b) ${\bf{D}}^{\pm}_{y}(k,l)$ $=\sum_{a=1}^{N_{x}}\sum_{b=1}^{N_{y}}\left(-\int_{\mathcal{T}_{a,b}}\partial_{y}\eta_{k}\eta_{l}d\mathbf{x}-\int_{x_{a-\frac{1}{2}}}^{x_{a+\frac{1}{2}}}\left(\eta_{k}(x,y_{b+\frac{1}{2}}^{+},y)-\eta_{k}(x,y_{b+\frac{1}{2}}^{-})\right)\eta_{l}(x,y_{b+\frac{1}{2}}^{\pm})dx\right),$
(30c) ${\mathbf{\Sigma}}_{s}(k,l)$ $=\sum_{a=1}^{N_{x}}\sum_{b=1}^{N_{y}}\int_{\mathcal{T}_{a,b}}\sigma_{s}\eta_{k}\eta_{l}d\mathbf{x},\qquad{\mathbf{\Sigma}}_{t}(k,l)=\sum_{a=1}^{N_{x}}\sum_{b=1}^{N_{y}}\int_{\mathcal{T}_{a,b}}\sigma_{t}\eta_{k}\eta_{l}d\mathbf{x}.$
The $k$-th element of vectors $\bm{G}$ and $\bm{g}_{j}^{\textrm{bc}}$ is $\bm{G}(k)=\sum_{a=1}^{N_{x}}\sum_{b=1}^{N_{y}}\int_{\mathcal{T}_{a,b}}G\eta_{k}d\mathbf{x}$ and
$\bm{g}_{j}^{\textrm{bc}}(k)=\sum_{b=1}^{N_{y}}\Big($ $-\max(\bm{\Omega}_{j,x},0)\int_{y_{b-\frac{1}{2}}}^{y_{b+\frac{1}{2}}}g(x_{L},y)\eta_{k}(x_{\frac{1}{2}}^{+},y)dy$
$+\min(\bm{\Omega}_{j,x},0)\int_{y_{b-\frac{1}{2}}}^{y_{b+\frac{1}{2}}}g(x_{R},y)\eta_{k}(x_{N_{x}+\frac{1}{2}}^{-},y)dy\Big)$
$+\sum_{a=1}^{N_{x}}\Big($ $-\max(\bm{\Omega}_{j,y},0)\int_{x_{a-\frac{1}{2}}}^{x_{a+\frac{1}{2}}}g(x,y_{B})\eta_{k}(x,y_{\frac{1}{2}}^{+})dx$
(31) $+\min(\bm{\Omega}_{j,y},0)\int_{x_{a-\frac{1}{2}}}^{x_{a+\frac{1}{2}}}g(x,y_{T})\eta_{k}(x,y_{N_{y}+\frac{1}{2}}^{-})dx\Big).$